# Gauge fields induced by curved spacetime

### v1

@@ Abstract
where $\gamma^{\mu}$ are the Dirac gamma matrices satisfying $\{\gamma^{\mu},\gamma^{\nu}\}=2\eta_{\mu\nu}$ with $\eta_{\mu\nu}=\operatorname{diag}{(1,-1)}$ the Minkowski metric, $\sqrt{-g}$ the square root of the determinant of the metric, and the zweibein $g_{\mu\nu}=e^{a}{}_{\mu}e^{b}{}_{\nu}\eta_{ab}$, $g^{\mu\nu}=e_{a}{}^{\mu}e_{b}{}^{\nu}\eta^{ab}$, with $g^{\mu\nu}=(g_{\mu\nu})^{-1}$. For two-component spinors, one can choose any combination of Pauli matrices, e.g., $\gamma^{0}=\sigma_{0}$ and $\gamma^{1}=\mathrm{i}\sigma_{x}$, satisfying the gamma-matrix algebra. Consider the metric
@@
which describes the time evolution of the spinor field, with $H$ the Hamiltonian density. Note that the Hamiltonian $H$ is not hermitian for $\partial_{1}\alpha(x)\neq 0$, since $(\mathrm{i}\gamma_{0}\gamma^{1})^{\dagger}=-\mathrm{i}\gamma_{0}\gamma^{1}$. The corresponding Hamiltonian can be regularized on a discrete lattice using

### v2

@@ Abstract
where $\psi$ is a spinor, $\gamma^{\mu}$ the flat spacetime Dirac gamma matrices satisfying $\{\gamma^{\mu},\gamma^{\nu}\}=2\eta^{\mu\nu}$ with $\eta_{\mu\nu}=\operatorname{diag}{(1,-1)}$ the Minkowski metric, $\sqrt{-g}$ the square root of the determinant of the metric, and the zweibein $g_{\mu\nu}=e^{a}{}_{\mu}e^{b}{}_{\nu}\eta_{ab}$, and $\eta_{ab}=e_{a}{}^{\mu}e_{b}{}^{\nu}g_{\mu\nu}$, with $g^{\mu\nu}=(g_{\mu\nu})^{-1}$. In the Weyl representation $\gamma^{0}=\sigma_{x}$ and $\gamma^{1}=\mathrm{i}\sigma_{y}$. Consider the metric
@@
which describes the time evolution of the spinor field, with $H$ the Hamiltonian density. The corresponding Hamiltonian can be regularized on a discrete lattice using